How much energy is absorbed when 6.0 g of water boils?

Given:

Mass of water 6.0 g

Molar enthalpy of fusion (ΔHf) of ice is 6.009 kJ/mol

Molar enthalpy of vaporization (ΔHv) of water is 40.79 kJ/mol

Equation: q=mol*ΔHv
1
Expert's answer
2020-05-01T14:41:48-0400

q = mol*ΔHv

n(H2O) = 6/18 = 0.3333 mol

q = 0.3333*40.79 = 13.6 kJ